# Structure of the Cayley-Dickson algebras

### v1

@@ 2 Graded Cayley-Dickson Construction
Anticommutation is proved by induction starting with Lemma 3 for $o_{1}o_{2}=-o_{2}o_{1}$. This is the $n=2$ level but is provided as the basis of induction in the proof of the lemma. By the symmetry of the construction, this result applies at all levels $n>2$ for these pure elements. Assuming the $n-1$ product is true, $a_{\alpha 0}b_{\beta 0}=-b_{\beta 0}a_{\alpha 0}$, for $\alpha,\beta\in{\mathrm{B}(n-1)}$, not all zeros, and $\alpha\neq\beta$, then multiplying by $o_{n}$ produces a pyramid whereby the left-hand half does not contain $o_{n}$ and the right-hand half reflects this structure but includes $o_{n}$ for all elements. Proving $xy=-yx$ for the $o_{n}$ case is divided into three equations which need to be proved; $o_{n}a_{\alpha 0}=-a_{\alpha 0}o_{n}$, $a_{\alpha 1}b_{\beta 0}=-b_{\beta 0}a_{\alpha 1}$ and $a_{\alpha 1}b_{\beta 1}=-b_{\beta 1}a_{\alpha 1}$, for $\alpha,\beta$ not all zeros. Lemma 3 has already shown the first case and found that $a^{*}_{\alpha}=-a_{\alpha}$ since $\alpha$ is not all zeros. Hence, the second case with $\beta$ not all zeros, is
$\begin{split}{a}_{\alpha 1}{b}_{\beta 0}&=(0,a_{\alpha 1})(b_{\beta 0},0)=(0,a_{\alpha 1}b_{\beta 0}^{*})\text{ and}\\ {b}_{\beta 0}{a}_{\alpha 1}&=(b_{\beta 0},0)(0,a_{\alpha 1})=(0,a_{\alpha 1}b_{\beta 0})=-{a}_{\alpha 1}b_{\beta 0}\text{.}\end{split}$
@@
$\begin{split}{a}_{\alpha 1}{b}_{\beta 1}&=(0,a_{\alpha 1})(0,b_{\beta 1})=(-b_{\beta 1}^{*}a_{\alpha 1},0)\text{ and}\\ {b}_{\beta 1}{a}_{\alpha 1}&=(0,b_{\beta 1})(0,a_{\alpha 1})=(-a_{\alpha 1}^{*}b_{\beta 1},0)\text{.}\end{split}$ (7)
These anticommutator terms occur in the same position on the right-hand half as the terms without $o_{n}$ on the left-hand half of the pyramid. After the first pyramid multiplication they have identical paths through the each side of the pyramid apart from the minus signs and the conjugate. The conjugate descends the path until the first 1 in $\alpha$ from the right and changes the sign and disappears for the rest of the path, $a_{\alpha 1}^{*}=-a_{\alpha 1}$ and similarly, $a_{\beta 1}^{*}=-a_{\beta 1}$. Hence the minus signs are cancelled in (7) and so using symmetry and arguing by induction with $a_{\alpha 0}b_{\beta 0}-b_{\beta 0}a_{\alpha 0}=0$ true at level 2 and assumed true at level $n-1$, then
$-b_{\beta 1}^{*}a_{\alpha 1}-a_{\alpha 1}^{*}b_{\beta 1}=a_{\alpha 0}b_{\beta 0}+b_{\beta 0}a_{\alpha 0}=0\text{.}$
@@
For $x$ and $y$ distinct blades with $\epsilon_{i}=1$, $i\in{\mathbb{N}}_{1}^{n}$, then $yxx=-y$ and $x(xy)=-y$.
@@
The later statement is true if the former is true, $x(xy)=-(xy)x=yxx$, if $y\neq x$ and obviously $x(xx)=-x$ otherwise. The proof of the former proceeds via induction. It works for ${\mathbb{H}}$ which needs to be checked for the six combinations of two from three elements using the cyclic product rules (8),
@@
This leaves $xo_{n}o_{n}$ where $\beta$ is all zeros in the last two equations. In that case the conjugate has no effect after the first step and $-b_{\beta 1}^{*}(b_{\beta 1}x)=-b_{\beta 1}(b_{\beta 1}^{*}x)=-x$. Hence ${a}_{\alpha}{b}_{\beta}{b}_{\beta}=-{a}_{\alpha}$ for all $\alpha,\beta\in{\mathrm{B}(n)}$. Note that while $xyzz=-xy$ it is not true that $x(xyz)=-yz$ for a distinct blade $z$. For example, $o_{1}(o_{1}o_{2}o_{3})=o_{23}$. ∎
@@
All of the triples in (LABEL:eqn:qa7) are quaternion-like, either isomorphic or anti-isomorphic to quaternions, with cyclic product rules and the first one is the quaternion subset from level 2. The next two have the same structure and would be called quaternions if the basis where reordered. The last triple provides rotations between the three generators, which is a feature of quaternions if the generators represent three dimensional vector space, ${\mathbb{R}}^{3}$. The three remaining triples can not be interpreted in this way but the graded product of the three blades in all seven triples is seen to be a scalar, either ±$\pm 1$. The remaining triples will be shown to be non-associative. For example, Lemma 2 has $o_{1}o_{2}o_{3}=o_{123}$ whereas expanding the pyramid for $o_{1}(o_{2}o_{3})$ gives the term $-a_{100}b_{011}$ in the $o_{123}$ position so that $o_{1}(o_{2}o_{3})=-o_{123}$, hence the associator is non-zero and this triple is non-associative, $[o_{1},o_{2},o_{3}]=2o_{123}$.
@@
The Cayley-Dickson series for power-associative algebras after level $3$ with $\epsilon_{i}=1$ for $i\in{\mathbb{N}}_{1}^{n}$ is designated ${\mathbb{U}_{m}}$, $m=n-3$, which refers to the ultronion algebras. Thus the normed algebras, ${\mathbb{R}}$, ${\mathbb{C}}$, ${\mathbb{H}}$ and ${\mathbb{O}}$, are not included in the ${\mathbb{U}_{m}}$ series and the series starts with the sedenions, ${\mathbb{S}}\equiv{\mathbb{U}_{1}}$. The split-algebras are denoted ${\rm\widetilde{\mathbb{U}}_{m}}$ and later it is shown they are isomorphic to each other, for any dimension, for any positive number of $\epsilon_{i}=-1$, $i\in{\mathbb{N}}_{1}^{n}$. The proof of uniqueness of ${\rm\widetilde{\mathbb{U}}_{n}}$ also applies to split versions of ${\mathbb{C}}$ (Lorentz numbers), ${\mathbb{H}}$ and ${\mathbb{O}}$.
@@ 3 Structure Theorems
There are $N^{3}$ scaled triple products $bcd$ and $\binom{N}{3}$ naturally ordered triples with blades $b<c<d$ where $N=2^{n}-1$ and $n$ is the number of generators. The ordered triples are the same as removing all pairs and triples of the same element from the $N^{3}$ set and dividing by the number of permutations, $6$. A triad is defined as ordered triples containing distinct blades whereby $b<c<d$ so that $bcd$ is non-zero. Hence basis element $a=bcd$ can be defined for triads and $a$ is distinct from $b,c,d$, which follows from Lemma 1. Of course, there are 5 unordered triples for each triad.
@@
The $24$ permutations of triad products generate three types of associativity, called Types 1, 2 and 3. These are shown in Table 1 where equivalence of associators means both are zero or both are non-zero, noting that only 12 permutations for ordered $b,c,d$ and $a=bcd$ are required because $[a,b,c]=-[c,b,a]$.
@@
$[b,a,c]\approx[b,d,c]\approx[a,b,d]\approx[a,c,d]$ Type 1
$[a,b,c]\approx[b,c,d]\approx[b,a,d]\approx[a,d,c]$ Type 2
$[a,c,b]\approx[c,b,d]\approx[a,d,b]\approx[c,a,d]$ Type 3
@@
Excluding complete associativity for triads where all associativity types are zero means that $a(bc)=-(bc)a=cba$ if $a$ is not a scaling of $bc$ and similarly for $abc=(ab)c=c(ba)$. Hence $[a,b,c]=abc-a(bc)=c(ba)-(cb)a=-[c,b,a]$. This is also true for associative triple products, where $a$ is a scalar, because both sides are zero. Now, if $b,c,d$ are not associative, the relation $a=bcd$ means $ad^{-1}=bc$ or $d=-a(bc)^{-1}=-bca/b^{2}/c^{2}$, noting that $(bc)^{2}=-b^{2}c^{2}$. First, Type 2 and Type 3 associativity are analysed.
Assuming Type 2 associativity $[a,b,c]=abc-a(bc)=0$ or $[c,b,a]=cba-c(ba)=0$, then substituting $d$ into the other three associators in the second row in Table 1, scaled by $(bc)^{2}$, gives
@@
Assuming Type 3 associativity $[a,c,b]=acb-a(cb)=0$ or $[b,c,a]=bca-b(ca)=0$, then substituting $d$ into the other three associators in the third row of Table 1, scaled by $(bc)^{2}$, gives
@@
Assuming Type 1 associativity $[b,a,c]=bac-b(ac)=0$ or $[c,a,b]=cab-c(ab)=0$, then substituting $d$ into the other three associators in the first row of Table 1, scaled by $(bc)^{2}$, gives
@@
If Type 3 associativity is also assumed with Type 1 then $bca=b(ca)$ or $cba=-b(ca)$ and, also again using $[c,a,b]=0$, the three equations become
@@
Any two distinct blades of ${\mathbb{H}}$, ${\mathbb{O}}$ or ${\mathbb{U}_{m}}$, $m>0$, generate an algebra isomorphic to quaternions which is thus associative.
@@
From Lemma 1, multiplying any two distinct blades generates a third distinct blade. Since the generated blade is the symmetric difference then multiplying this third element by one of the first two will recover the other of the first two. All such products are closed under multiplication and forms a Cayley-Dickson algebra, by Theorem 1, because distinct blades anti-commute and have negative square in this case. For any two blades, $b$ and $c$ then $bc=d$, $cd=c(bc)=-bcc=b$ and $db=-cbb=c$, by Lemma 4. These are the cyclic, quaternion-like product rules and thus the triad $b$, $c$ and $bc$ is isomorphic to ${\mathbb{H}}$. The triple associator shows associativity for the triad,
$\begin{split}T(b,c,bc)&=b(bc)c-b(bcc)-bc(cb)+bc(cb)+cb(bc)-c(b(bc))\\ &=-bcbc+bb-bc(bc)+c(bcb)\\ &=+1-1+1-1\\ &=0\text{.}\\[-17.22217pt] \end{split}$
@@
Any ultacomplex algebra, ${\mathbb{U}_{m}}$, $m>0$, can be generated by $n=m+3$ distinct blades, $o_{\alpha_{k}},\forall\>k\in{\mathbb{N}}_{1}^{n}$, with indices $\alpha_{k}$ that include all $n$ indices an odd number of times over all $k$.
@@
Combinations of $b,c,d$ give seven distinct blades by Lemma 1, $bc,bd,cd,bcd$, which are isomorphic to octonions because they satisfy the rules of ${\mathbb{U}_{n}}$ and these are the same as ${\mathbb{O}}$ by Theorem 1. Any $n$ such terms similarly generate $12^{n}-$ 1 distinct terms if, by Lemma 1, the symmetric difference of their grades do not become the empty set, which is guaranteed by the condition that each generator is represented an odd number of times in the indices. Hence they are isomorphic to ${\mathbb{U}_{n-3}}$, by Theorem 1. ∎
@@
which contradicts Type 3 associativity. This proves existence of associativity pairs for non-associativity. For necessity, if any associator is non-zero then since $a=bcd$ then the associator must be ±$\pm 2a$. If only one associator type is zero and the other two are non-zero then from (10), the triple associator becomes
@@ 0 $\star$ 0 Type B
The simplified identity means $[bc,d,b]\neq 0$ which implies A or X non-associativity if $bc$ is interpreted as a new $c$. Malcev algebras define non-associative Lie algebras and this is analysed later where B and C non-associativity types are also found. For quaternion-like triads, $d=bc$ can be substituted into (11),
$bcdb+bc(db)=bc(bc)b+bc(bcb)=-b+bcc=-2b$
@@
1 $d(b(dc))-dbdc=0$ B or X
2 $b(d(cd))-bdcd=0$ C or X
3 $db(cd)-d(bc)d=0$ B or X if Malcev,
@@
Substituting the first equation into Type 2 associativity is $[b,c,d]=bcd-b(cd)=2bcd$, which is non-zero. Substituting the second into Type 3 associativity is $[c,b,d]=cbd-c(bd)=2cbd$, which is also non-zero. Hence Moufang loop non-associativity of Type 1 and 2 corresponds to B-non-associativity and C-non-associativity, respectively, or X-non-associativity in both cases.
@@
so $cd=bc(db)$. Applying (11) so that $cd=-bcdb$ and substituting into Type 2 associativity gives,
$[b,c,d]=bcd-b(cd)=bcd+b(bcdb)=bcd-bcdbb=bcd+bcd=2bcd\neq 0\text{.}$
This implies type B or X non-associativity. If Mal’cev’s identity is not satisfied then (11) takes the opposite sign, $[bc,d,b]=0$, and Type 2 associativity is then zero so that A or C non-associativity holds. This completes the proof and it is noted that it is possible to use combinations of the Moufang types along with Mal’cev’s identity to distinguish all four types of non-associativity A, B, C and X. ∎
@@
All these octonion triads have X-non-associativity and it will soon be proved that the cycles must also be of type X. Remember $[o_{1},o_{2},o_{3}]=2o_{123}$ and multiplying terms by distinct basis elements does not change the non-associativity type. To achieve this we need another generator, $o_{4}$, for example, so that for $b=o_{1}$, $c=o_{2}$ and $d=o_{34}$ then Type 2 is associative, $[o_{2},o_{3},o_{34}]=-o_{1234}+o_{1234}=0$ and Type 1 is non-associative, $[o_{1},o_{34},o_{2}]=-o_{1234}-o_{1234}\neq 0$.

### v2

@@ 2 Graded Cayley-Dickson Construction
Anticommutation is proved by induction starting with Lemma 3 for $o_{1}o_{2}=-o_{2}o_{1}$. This is the $n=2$ level but is provided as the basis of induction in the proof of the lemma. By the symmetry of the construction, this result applies at all levels $n>2$ for these pure elements. Assuming the n1n-1 product is true, $a_{\alpha 0}b_{\beta 0}=-b_{\beta 0}a_{\alpha 0}$, for $\alpha,\beta\in{\mathrm{B}(n-1)}$, not all zeros, and $\alpha\neq\beta$, then multiplying by $o_{n}$ produces a pyramid whereby the left-hand half does not contain $o_{n}$ and the right-hand half reflects this structure but includes $o_{n}$ for all elements. Proving xy=yxxy=-yx for the $o_{n}$ case is divided into three equations which need to be proved; $o_{n}a_{\alpha 0}=-a_{\alpha 0}o_{n}$, $a_{\alpha 1}b_{\beta 0}=-b_{\beta 0}a_{\alpha 1}$ and $a_{\alpha 1}b_{\beta 1}=-b_{\beta 1}a_{\alpha 1}$, for $\alpha,\beta$ not all zeros. Lemma 3 has already shown the first case and found that $a^{*}_{\alpha}=-a_{\alpha}$ since $\alpha$ is not all zeros. Hence, the second case with $\beta$ not all zeros, is
[…]
@@
[…]
These anticommutator terms occur in the same position on the right-hand half as the terms without $o_{n}$ on the left-hand half of the pyramid. After the first pyramid multiplication they have identical paths through the each side of the pyramid apart from the minus signs and the conjugate. The conjugate descends the path until the first 1 in $\alpha$ from the right and changes the sign and disappears for the rest of the path, $a_{\alpha 1}^{*}=-a_{\alpha 1}$ and similarly, $a_{\beta 1}^{*}=-a_{\beta 1}$. Hence the minus signs are cancelled in (8) and so using symmetry and arguing by induction with $a_{\alpha 0}b_{\beta 0}-b_{\beta 0}a_{\alpha 0}=0$ true at level 2 and assumed true at level n1n-1, then
$b_{\beta 1}^{*}a_{\alpha 1}-a_{\alpha 1}^{*}b_{\beta 1}=a_{\alpha 0}b_{\beta 0}+b_{\beta 0}a_{\alpha 0}=0\text{.}$
@@
For $x$ and $y$ distinct blades with $\epsilon_{i}=1$, $i\in{\mathbb{N}}_{1}^{n}$, then yxx=yyxx=-y and x(xy)=yx(xy)=-y.
@@
The later statement is true if the former is true, x(xy)=(xy)x=yxxx(xy)=-(xy)x=yxx, if $y\neq x$ and obviously x(xx)=xx(xx)=-x otherwise. The proof of the former proceeds via induction. It works for $\mathbb{H}$} which needs to be checked for the six combinations of two from three elements using the cyclic product rules (9),
@@
This leaves $xo_{n}o_{n}$ where $\beta$ is all zeros in the last two equations. In that case the conjugate has no effect after the first step and $b_{\beta 1}^{*}(b_{\beta 1}x)=-b_{\beta 1}(b_{\beta 1}^{*}x)=-x$. Hence ${a}_{\alpha}{b}_{\beta}{b}_{\beta}=-{a}_{\alpha}$ for all $\alpha,\beta\in{\mathrm{B}(n)}$. Note that while xyzz=xyxyzz=-xy, it is not necessarily true that x(xyz)=yzx(xyz)=-yz for distinct blades $x,y,z$. For example, $o_{1}(o_{1}o_{2}o_{3})=o_{23}$. ∎
@@
All of the triples in (LABEL:eqn:qa7) are quaternion-like, either isomorphic or anti-isomorphic with quaternions with cyclic product rules, and the first one is the quaternion subset from level 2. The next two have the same structure and would be called quaternions if the basis where reordered. The last triple provides rotations between the three generators, which is a feature of quaternions if the generators represent three dimensional vector space, ${\mathbb{R}}^{3}$. The three remaining triples can not be interpreted in this way but the graded product of the three blades in all seven triples is seen to be a scalar, either $\pm 1$. The remaining triples will be shown to be non-associative. For example, Lemma 2 has $o_{1}o_{2}o_{3}=o_{123}$ whereas expanding the pyramid for $o_{1}(o_{2}o_{3})$ gives the term $a_{100}b_{011}$ in the $o_{123}$ position so that $o_{1}(o_{2}o_{3})=-o_{123}$, hence the associator is non-zero and this triple is non-associative, $o_{1},o_{2},o_{3}]=2o_{123}$.
@@
The Cayley-Dickson series for power-associative algebras after $A_{3}$, with $\epsilon_{i}=1$ for $i\in{\mathbb{N}}_{1}^{n}$, are designated $\mathbb{U}_{m}$}, m=n3m=n-3, which refers to the ultronion algebras. Thus the normed algebras, $\mathbb{R}$}, $\mathbb{C}$}, $\mathbb{H}$} and $\mathbb{O}$}, are not included in the $\mathbb{U}_{m}$} series and the series starts with the sedenions, ${\mathbb{S}}\equiv{\mathbb{U}_{1}}$. The split algebras are denoted $\rm\widetilde{\mathbb{U}}_{m}$} and later it is shown they are isomorphic to each other, for any dimension, for any positive number of $\epsilon_{i}=-1$, $i\in{\mathbb{N}}_{1}^{n}$. The proof of uniqueness of $\rm\widetilde{\mathbb{U}}_{n}$} also applies to split versions of $\mathbb{C}$} (Lorentz numbers), $\mathbb{H}$} and $\mathbb{O}$}.
@@ 3 Structure Theorems
There are $N^{3}$ scaled triple products bcdbcd and $\binom{N}{3}$ naturally ordered triples with blades $b<c<d$ where $N=2^{n}-1$ and $n$ is the number of generators. The ordered triples are the same as removing all pairs and triples of the same element from the $N^{3}$ set and dividing by the number of permutations, $6$. A triad is defined as ordered triples containing distinct blades whereby $b<c<d$ so that bcdbcd is non-zero. Hence basis element $a=bcd$ can be defined for triads and $a$ is distinct from $b,c,d$, which follows from Lemma 1. Of course, there are 5 unordered triples for each triad.
@@
The $24$ permutations of triad products generate three types of associativity, called Types 1, 2 and 3. These are shown in Table 1 where equivalence of associators means both are zero or both are non-zero, noting that only 12 permutations for ordered $b,c,d$ and $a=bcd$ are required because [a,b,c]=[c,b,a][a,b,c]=-[c,b,a].
@@
$b,a,c]\approx[b,d,c]\approx[a,b,d]\approx[a,c,d]$ Type 1
$a,b,c]\approx[b,c,d]\approx[b,a,d]\approx[a,d,c]$ Type 2
$a,c,b]\approx[c,b,d]\approx[a,d,b]\approx[c,a,d]$ Type 3
@@
Excluding complete associativity for triads where all associativity types are zero means that a(bc)=(bc)a=cbaa(bc)=-(bc)a=cba if $a$ is not a scaling of $bc$ and similarly for $abc=(ab)c=c(ba)$. Hence [a,b,c]=abca(bc)=c(ba)(cb)a=[c,b,a][a,b,c]=abc-a(bc)=c(ba)-(cb)a=-[c,b,a]. This is also true for associative triple products, where $a$ is a scalar, because both sides are zero. Now, if $b,c,d$ are not associative, the relation $a=bcd$ means $ad^{-1}=bc$ or $d=-a(bc)^{-1}=-bca/b^{2}/c^{2}$, noting that $(bc)^{2}=-b^{2}c^{2}$. First, Type 2 and Type 3 associativity are analysed.
Assuming Type 2 associativity, [a,b,c]=abca(bc)=0[a,b,c]=abc-a(bc)=0 or [c,b,a]=cbac(ba)=0[c,b,a]=cba-c(ba)=0, then substituting $d$ into the other three associators in the second row in Table 1, scaled by $(bc)^{2}$, gives
@@
Assuming Type 3 associativity, [a,c,b]=acba(cb)=0[a,c,b]=acb-a(cb)=0 or [b,c,a]=bcab(ca)=0[b,c,a]=bca-b(ca)=0, then substituting $d$ into the other three associators in the third row of Table 1, scaled by $(bc)^{2}$, gives
@@
Assuming Type 1 associativity, [b,a,c]=bacb(ac)=0[b,a,c]=bac-b(ac)=0 or [c,a,b]=cabc(ab)=0[c,a,b]=cab-c(ab)=0, then substituting $d$ into the other three associators in the first row of Table 1, scaled by $(bc)^{2}$, gives
@@
If Type 3 associativity is also assumed with Type 1 then $bca=b(ca)$ or cba=b(ca)cba=-b(ca) and, also again using $[c,a,b]=0$, the three equations become
@@
Any two distinct blades of $\mathbb{H}$}, $\mathbb{O}$} or $\mathbb{U}_{m}$}, $m>0$, generate an algebra isomorphic or anti-isomorphic to quaternions which is thus associative.
@@
From Lemma 1, multiplying any two distinct blades generates a third distinct blade. Since the generated blade is the symmetric difference then multiplying this third element by one of the first two will recover the other of the first two. All such products are closed under multiplication and forms a Cayley-Dickson algebra, by Theorem 1, because distinct blades anti-commute and have negative square in this case. For any two blades, $b$ and $c$ then $bc=d$, cd=c(bc)=bcc=bcd=c(bc)=-bcc=b and db=cbb=cdb=-cbb=c, by Lemma 4. These are the cyclic, quaternion-like product rules and thus the triad $b$, $c$ and $bc$ is isomorphic to $\mathbb{H}$}. The triple associator shows associativity for the triad,
[…]
@@
Any ultacomplex algebra, $\mathbb{U}_{m}$}, $m>0$, as well as $\mathbb{O}$}, can be generated by $n=m+3$ distinct blades, $o_{\alpha_{k}},\forall\>k\in{\mathbb{N}}_{1}^{n}$, with indices $\alpha_{k}$ that include all $n$ indices an odd number of times over all $k$. For convenience, allow $m=0$ as ${\mathbb{U}_{0}}={\mathbb{O}}$.
@@
Combinations of $b,c,d$ give seven distinct blades by Lemma 1, $bc,bd,cd,bcd$, which are isomorphic to octonions because they satisfy the rules of $\mathbb{U}_{n}$} and these are the same as $\mathbb{O}$} by Theorem 1. Any $n$ such terms similarly generate $2^{n}-1$ distinct terms if, by Lemma 1, the symmetric difference of their grades do not become the empty set, which is guaranteed by the condition that each generator is represented an odd number of times in the indices. Hence they are isomorphic to $\mathbb{U}_{n-3}$}, by Theorem 1. ∎
@@
which contradicts Type 3 associativity. This proves existence of associativity pairs for non-associativity. For necessity, if any associator is non-zero then since $a=bcd$ then the associator must be $\pm 2a$. If only one associator type is zero and the other two are non-zero then from (11), the triple associator becomes
@@ 0 $\star$ 0 Type B
The simplified identity means $bc,d,b]\neq 0$ which implies A or X non-associativity if $bc$ is interpreted as a new $c$. Malcev algebras define non-associative Lie algebras and this is analysed later where B and C non-associativity types are also encountered. For quaternion-like triads, $d=bc$ can be substituted into (12),
bcdb+bc(db)=bc(bc)b+bc(bcb)=b+bcc=2bbcdb+bc(db)=bc(bc)b+bc(bcb)=-b+bcc=-2b
@@
1 d(b(dc))dbdc=0d(b(dc))-dbdc=0 B or X
2 b(d(cd))bdcd=0b(d(cd))-bdcd=0 C or X
3 db(cd)d(bc)d=0db(cd)-d(bc)d=0 B or X if Malcev,
@@
Substituting the first equation into Type 2 associativity is [b,c,d]=bcdb(cd)=2bcd[b,c,d]=bcd-b(cd)=2bcd, which is non-zero. Substituting the second into Type 3 associativity is [c,b,d]=cbdc(bd)=2cbd[c,b,d]=cbd-c(bd)=2cbd, which is also non-zero. Hence Moufang loop non-associativity of Type 1 and 2 corresponds to B-non-associativity and C-non-associativity, respectively, or X-non-associativity in both cases.
@@
so $cd=bc(db)$. Applying (12) so that cd=bcdbcd=-bcdb and substituting into Type 2 associativity gives,
$b,c,d]=bcd-b(cd)=bcd+b(bcdb)=bcd-bcdbb=bcd+bcd=2bcd\neq 0\text{.}$
This implies type B or X non-associativity. If Mal’cev’s identity is not satisfied then (12) takes the opposite sign, $[bc,d,b]=0$, and Type 2 associativity is then zero so that A or C non-associativity holds. This completes the proof and it is noted that it is possible to use combinations of the Moufang types along with Mal’cev’s identity to distinguish all four types of non-associativity A, B, C and X.
Note that some authors[22, 23] include a fourth Moufang type, $db(cd)=d(bcd)$. But this is the same as the third type since d(bcd)=bcdd=d(bc)dd(bcd)=-bcdd=d(bc)d. ∎
@@
All these octonion triads have X-non-associativity and it will soon be proved that the cycles must also be of type X. Remember $o_{1},o_{2},o_{3}]=2o_{123}$ and multiplying terms by distinct basis elements does not change the non-associativity type. To achieve this we need another generator, $o_{4}$, for example, so that for $b=o_{1}$, $c=o_{2}$ and $d=o_{34}$ then Type 2 is associative, $o_{2},o_{3},o_{34}]=-o_{1234}+o_{1234}=0$ and Type 1 is non-associative, $o_{1},o_{34},o_{2}]=-o_{1234}-o_{1234}\neq 0$.